OpenStudy (anonymous):

How many liters of oxygen gas, at standard temperature and pressure, will react with 35.4 grams of calcium metal? Show all of the work used to solve this problem. 2 Ca + O2 yields 2 CaO

OpenStudy (anonymous):

2mol Ca react with 1 mol O2 Molar mass Ca = 40.08g/mol 35.4g = 35.4/40.08 = 0.884 mol Ca This will react with 0.442 mol O2 At STP, 1mol O2 = 22.4 litres 0.442 mol O2 = 22.4*0.442 = 9.9 litres.

OpenStudy (anonymous):

No.. This is not a practice exam. This is a module quiz from FLVS and it's cheating if you use this answer. Plus it has a obvious mistake in it. Mr. Carlyle FLVS Chemistry Instructor This question is a violation of the OpenStudy Guidelines. The question is from an online school plus a violation by solicitation for members to unknowingly assist them cheating on an exam. 30653065chem30653065